Molality of iron(II) chloride in a solution is 2.736 mol·kg-1.

What is the mole fraction of FeCl2 in the solution?

M(FeCl2) =126.75 g·mol-1M(H2O) =18.02 g·mol-1


Present your numerical answer to 3 significant figures 

1
Expert's answer
2021-01-15T06:38:17-0500

mole fraction x(FeCl2) = n(FeCl2)/(n(H2O) + n(FeCl2)) = 2.736/(2.736+1000/18.02) = 0.0470